# dMel: Speech Tokenization made Simple

### v1

@@ Abstract
Large language models have revolutionized natural language processing by leveraging self-supervised pretraining on vast textual data. Inspired by this success, researchers have investigated complicated speech tokenization methods to discretize continuous speech signals so that language modeling techniques can be applied to speech data. However, existing approaches either model semantic tokens, potentially losing acoustic information, or model acoustic tokens, risking the loss of semantic information. Having multiple token types also complicates the architecture and requires additional pretraining. Here we show that discretizing mel-filterbank channels into discrete intensity bins produces a simple representation (dMel), that performs better than other existing speech tokenization methods. Using a transformer decoder-only architecture for speech-text modeling, we comprehensively evaluate different speech tokenization methods on speech recognition (ASR) and speech synthesis (TTS). Our results demonstrate the effectiveness of dMel in achieving high performance on both tasks within a unified framework, paving the way for efficient and effective joint modeling of speech and text. Our code will be published soon.
@@ 1 Introduction
Current speech tokenization approaches can be broadly categorized into two types: semantic tokens and acoustic tokens. Semantic tokens, extracted from self-supervised (SSL) pretrained speech models [2, 14], where the speech signal is first encoded into speech representations and then clustered into semantic tokens with $k$-means method. However, such SSL pretrained models are not useful for high fidelity speech synthesis as speaker identity and other details of raw speech are lost in training [7]. Conversely, acoustic tokens can be obtained from audio compression models that are trained to compress the speech signal into codebook indices with residual vector quantization (RVQ) and reconstruction objectives [37, 11]. These tokens prioritize acoustic reconstruction but lose semantic information which can lead to poorer results in generating audio [31].
To combine the advantages of both semantic and acoustic tokens, AudioLM [7] proposed to model both semantic tokens and acoustic tokens with 3 stages: semantic modeling, coarse acoustic modeling, and fine acoustic modeling. The coarse-to-fine modeling strategy is designed to match the residual structure of RVQ based acoustic tokens. This solution addresses both content and speech quality, but its multi-stage hierarchical structure complicates the model and can lead to slower training and inference. Another solution is to combine the semantic and acoustic features together. Zhang et al. [40] proposed to distill the semantic tokens into the acoustic token’s first residual channel during the training of the RVQ model in a teacher-student manner. In this way, they showed that the new feature can preserve the semantic information better and also reconstruct high quality speech signals.
In this paper, we raise the following fundamental question – do we really need to separate speech into semantic and acoustic tokens first, and process them with idiosyncratic architectures ? We propose a simple alternative called dMel (see Figure 1) that discretizes mel-filterbanks energies directly into ordinal bins. Intriguingly, we find that discretizing Mel spectrograms has little impact on the ability of off-the-shelf mel-filterbank vocoders to reconstruct waveforms. [^1] We used vocoders from https://github.com/kan-bayashi/ParallelWaveGAN.
@@
WER of Reconstruction
Ground-truth 2.02
Parallel-WavGAN HifiGAN
Mel 2.13 2.08
dMel 2.23 2.11
In Table 1 we use two different vocoders to reconstruct waveforms from mel-filterbanks (Mel) and discretized mel-filterbanks (dMel) computed on them. We find that the word error rate (WER) of an ASR system run on the reconstructed waveforms, is quite similar to the WER of the same system run on the ground-truth audio showing that discretizing Mel has limited impact on information content.
By operating on the mel-filterbanks and preserving the frequency and intensity information (with some loss of resolution from discretization), dMel inherently preserves both semantic and acoustic information in a unified representation, without the need for separate tokenization or additional pretraining of a tokenization model. There are many advantages to discretizing mel-filterbanks:
  • Mel-filterbanks is an interpretable representation of speech signal, where both the semantic and acoustic information is preserved and discretization has little impact.
@@
  • Different channels of dMel do not have the complex hierarchical dependencies on each other that is typical of coarse-to-fine acoustic tokens; we find that they can be modeled independently in each frame using a simple decoder-only transformer architecture.
Through comprehensive evaluations, we show that using dMel allows us to use a single decoder-only model, and achieve high performance on both ASR and TTS tasks. The ASR task validates that dMel preserves semantic information, while the TTS task shows that dMel are useful for high-fidelity acoustic reconstruction of speech. We also compare dMel to other tokenization methods and find that dMel achieves the best WER for ASR task, which indicates that the semantic information is well preserved. Also, dMel achieves the lower WER score for TTS task when using WhisperX [4] for automatic evaluation and we find model trained with dMel can generate long and natural speech samples (See Supplementary Material).
Figure 1: Prior works on speech tokenization use either heavy self-supervised pretrained encoders [2, 14] to extract semantic tokens (and train a separate decoder for it) [19] or learn compression encoder-decoder models with residual vector quantizations [37, 11] to obtain acoustic tokens. By contrast we eliminate the encoder and simply discretize mel-filerbanks (dMel) to encode audio, and use a simple mel-filterbank vocoder [36] to reconstruct speech signals.
@@ 2 Method
In this section, we first introduce our proposed dMel speech tokenization method, which discretizes mel-filterbanks energies directly into bins. We then describe our unified transformer decoder-only model for ASR and TTS tasks, which leverages dMel for speech tokenization. The model architecture is illustrated in Figure 2.
@@
Different from existing VQ-VAE [7, 40, 17, 37] based speech tokenizers, we propose a discretized mel-filterbanks based speech tokenizer. The outline of the discretization method is shown in Figure 1. Later in the paper, we show that this tokenizer allows the model to process the input speech signal efficiently and capture the relevant acoustic features for both ASR and TTS tasks.
We denote tensors as $\mathbf{X}$ while $\mathbf{X}_{i,...}$ denote the $(i,...)$-th component of tensor $\mathbf{X}$. First, the speech tokenizer takes the input speech signal $\mathbf{x}$ and computes the mel-filterbanks representation $\mathbf{M}$:
@@
where $\text{Mel}(\cdot)$ represents the function that computes the mel-filterbanks, $\mathbf{M}\in\mathbb{R}^{T\times N}$, $N$ is the number of filterbanks and $T$ is the number of frames in the spectrogram.
@@
To discretize the mel-filterbanks representation $\mathbf{M}$ into speech tokens, we adopt a codebook $\mathbf{C}$. In this paper, we apply a simple linear discretization, so that the codebook $\mathbf{C}\in\mathbb{R}^{2^{K}}$ and its values are evenly spaced in the range of the mel-filterbanks values:
@@
In practice, we compute the minimum $m$ and maximum $M$ values of mel-filterbanks across the entire dataset to define the codebook $\mathbf{C}$. Then we map a magnitude $\mathbf{M}_{t,i}$ of every frequency channel $i=1\dots N$ for the time frame $t=1\dots T$ into a bin index of the codebook $\mathbf{C}$ in the following way:
@@
where $\mathbf{S}\in\mathbf{B}^{T\times N}$ represents the discretized mel-filterbanks (dMel) with $\mathbf{B}=\{j|j=1,2,3,\dots 2^{K}\}$ and $\mathbf{S}_{t}\in\mathbf{B}^{N}$ being the $t$-th speech token. As the codebook $\mathbf{C}$ has $2^{K}$ distinct values and thus number of bins $|\mathbf{B}|=2^{K}$, each speech token is represented by $N\cdot K$ bits where every $K$ bits are used to represent one of $N$ frequency channels.
@@
To reconstruct the speech signal $\mathbf{x}$ from the speech tokens $\mathbf{S}$, we first transform bin indices back to the mel-filterbanks representation via the codebook $\mathbf{C}$:
@@
Then, we apply a vocoder [36] to transform reconstructed mel-filterbanks $\hat{\mathbf{M}}_{t,i}$ back into the time domain signal $\mathbf{x}$. The vocoder is trained independently and is not part of the transformer decoder-based model.
@@ 2.2 Unified Speech-Text Transformer Decoder
Figure 2: Unified Speech-Text Transformer Decoder.
Modeling speech and text sequences jointly is essential for a model to understand and generate both modalities. However, it is challenging to design a unified model that can handle both speech-to-text and text-to-speech effectively. In this work, we apply a unified transformer decoder-only model that takes speech and text tokens as input and generates the output tokens in the target sequence. The model is trained in end-to-end on a combined dataset of speech and text pairs, enabling it to learn the joint representations for ASR and TTS tasks. As we show in the rest of the paper, the crucial part for the joint model training is the proper speech tokenization which dMel provides.
@@
$\mathbf{E^{\prime}}_{t}$ $=$ $\text{Concatenate}([\text{Embed}(\mathbf{S}_{t,1}),\text{Embed}(\mathbf{S}_{t,2}),\ldots,\text{Embed}(\mathbf{S}_{t,N})])$ (6)
$\mathbf{E}_{t}$ $=$ $\text{Linear}(\mathbf{E^{\prime}}_{t}),$ (7)
where $\mathbf{E}_{t}\in\mathbb{R}^{D}$ is the speech token representation. Here, for every time frame $t$ a speech token $\mathbf{S}_{t}$ is processed in parallel and independently for every frequency channel $i$ by $\text{Embed}(\mathbf{S}_{t,i})$ mapping, and then embeddings of all frequency channels are stacked together to form one vector representation $\mathbf{E^{\prime}}_{t}$ for the frame $t$. Finally, the speech token embeddings $\mathbf{E}_{t}$ are fed into the decoder-only transformer models for further processing.
We also implemented other popular speech tokenizers including HuBERT-KM [19] and SpeechTokenizer [40] for comparison. The main difference among these speech tokenizers is the codebook size and codes dimension, which are shown in Table 2. For both HuBERT-KM and SpeechTokenizer the speech tokens are mapped via a learnable linear layer from their dimension to the text embedding dimension $D$ before feeding into the transformer decoder-only model.
Table 2: Comparison between different speech tokenizers: dMel (ours), HuBERT-LM and SpeechTokenizer. For dMel we use $N=80$ mel-filterbanks and $162^{K}=$ 16 values of the codebook $\mathbf{C}$.
@@
To properly model multi-speaker data, we also include speaker embeddings as input to the transformer decoder. The speaker embeddings are extracted from an independent dvector [30] model. [^2] We use a pretrained model “Speaker Encoder” from the YourTTS [9] repository https://github.com/Edresson/YourTTS. We use a learnable linear layer to map the speaker embeddings to the same dimension as the speech and text token embeddings $D$. The speaker representation is optional for ASR task, but required for TTS task. Hence, during the training, it is applied for text-to-speech and ignored for speech-to-text.
@@
To capture the relative distances between tokens in the input sequence we apply multiplicative relative positional embedding RoPE [29]. This allows the model to learn the positional relationships between speech tokens, text tokens, and speaker embeddings, enhancing its ability to generate coherent output sequences. For positional embeddings we do not distinguish between text, speech and speaker embeddings having global positions notation across all kind of tokens.
2.3 Robust Training
Compared to language modeling, audio frames are highly redundant with strong local correlations. This makes long form generation difficult for models due to exposure bias [6]. To mitigate exposure bias during training, we apply span-masking [25] to the speech token context, masking out multiple random spans of speech frames. The model is trained to predict the next token based on the masked context. This context-masking strategy helps the model learn to generate accurate speech tokens in the presence of missing information, improving its robustness and generalization. It forces the model to attend to the text rather than copying previously inferred speech tokens due to learnt correlations. We also find that span-masking text tokens improves the ASR task.
@@ 3 Experiments
In this section, we conduct ASR and TTS experiments using dMel tokens with a unified transformer decoder-only model. We evaluate the performance of our model mainly on the LibriSpeech dataset and compare it with state-of-the-art speech tokenization methods, ASR and TTS models.
@@ 3.1 Training Data
We use several open-sourced datasets with paired speech and text transcription to conduct experiments: i) LibriSpeech [22] dataset (CC BY 4.0) consists of English speech recordings (960h, 16kHz) from various speakers ($\sim$2k) and conditions; ii) LibriTTS [38] (CC BY 4.0) dataset derived from LibriSpeech improves over it with the proper sentence split, text normalization and keeping 24kHz. It has around 500h; iii) VCTK [34] contains 44 hours of English speech (108 speakers); iv) LJSpeech [15] dataset (public domain in USA) is a single speaker English audio recordings of 16kHz with read speech from LibriVox. [^3] https://librivox.org/pages/public-domain/ While LibriSpeech is used to train ASR, TTS and joint models, LibriTTS, VCTK and LJSpeech are used to train TTS model only.

### v2

@@ Abstract
Large language models have revolutionized natural language processing by leveraging self-supervised pretraining on vast textual data. Inspired by this success, researchers have investigated complicated speech tokenization methods to discretize continuous speech signals so that language modeling techniques can be applied to speech data. However, existing approaches either model semantic (content) tokens, potentially losing acoustic information, or model acoustic tokens, risking the loss of semantic (content) information. Having multiple token types also complicates the architecture and requires additional pretraining. Here we show that discretizing mel-filterbank channels into discrete intensity bins produces a simple representation (dMel), that performs better than other existing speech tokenization methods. Using an LM-style transformer architecture for speech-text modeling, we comprehensively evaluate different speech tokenization methods on speech recognition (ASR) and speech synthesis (TTS). Our results demonstrate the effectiveness of dMel in achieving high performance on both tasks within a unified framework, paving the way for efficient and effective joint modeling of speech and text. The code will be open sourced.
@@ 1 Introduction
Current speech tokenization approaches can be broadly categorized into two types: semantic (content) tokens and acoustic tokens [^1] We use a word ‘semantic’ with the meaning of ‘content’ to keep prior work notation [7].. Semantic tokens, extracted from self-supervised (SSL) pretrained speech models [2, 16], where the speech signal is first encoded into speech representations and then clustered into semantic tokens with $k$-means method. However, such SSL pretrained models are not useful for high fidelity speech synthesis as speaker identity and other details of raw speech are lost in training [7]. Conversely, acoustic tokens can be obtained from audio compression models that are trained to compress the speech signal into codebook indices with residual vector quantization (RVQ) and reconstruction objectives [42, 11]. These tokens prioritize acoustic reconstruction but lose semantic information which can lead to poorer results in generating audio [38].
To combine the advantages of both semantic and acoustic tokens, AudioLM [7] proposed to model both semantic tokens and acoustic tokens with 3 stages: semantic modeling, coarse acoustic modeling, and fine acoustic modeling. The coarse-to-fine modeling strategy is designed to match the residual structure of RVQ based acoustic tokens. This solution addresses both content and speech quality, but its multi-stage hierarchical structure complicates the model and can lead to slower training and inference. Another solution is to combine the semantic and acoustic features together. Zhang et al. [46] proposed to distill the semantic tokens into the acoustic token’s first residual channel during the training of the RVQ model in a teacher-student manner. In this way, the new feature can preserve the semantic information better and also reconstruct high quality speech signals.
In this paper, we raise the following fundamental question – do we really need to separate speech into semantic and acoustic tokens first, and process them with idiosyncratic architectures? We propose a simple alternative called dMel (see Figure 1) that discretizes log mel-filterbanks (Mel) energies directly into ordinal bins. Intriguingly, we find that discretizing Mel has little impact on the ability of off-the-shelf Mel vocoders to reconstruct waveforms [^2] We used vocoders from https://github.com/kan-bayashi/ParallelWaveGAN.. In Table 1 we show different vocoders to reconstruct waveforms from Mel and discretized Mel (dMel) computed on them, as well as ASR models trained on Mel and dMel.
@@
Reconstruction WER (%) Recognition WER (%)
P-WaveGAN1 HifiGAN2 Seq2seq3 CTC4
Ground-truth 2.02 -
Mel 2.13 2.08 2.4 2.1
dMel 2.23 2.11 2.5 2.1
* [41]1, [21]2, [12]3, [13]4 Configurations are detailed in Sec. 3.
We find that the word error rate (WER) of an ASR system run on the reconstructed waveforms, is quite similar to the WER of the same system run on the ground-truth audio, showing that dMel captures the acoustic information needed to reconstruct good waveforms. Similarly, we find that the WER of ASR models trained on Mel and dMel are similar, indicating that dMel are good at preserving semantic content that can easily be found by ASR models. This shows that discretizing Mel has limited impact on information content.
By operating on the log mel-filterbanks and preserving the frequency and intensity information (with some loss of resolution from discretization), dMel inherently preserves both semantic and acoustic information in a unified representation, without the need for separate tokenization or additional pretraining of a tokenization model. There are many advantages to discretizing log mel-filterbanks:
  • Log mel-filterbanks is an interpretable representation of speech signal, where both the semantic and acoustic information is preserved. As discretization has little impact, dMel inherits the properties.
@@
  • Different channels of dMel do not have the complex hierarchical dependencies on each other that is typical of coarse-to-fine acoustic tokens; we find that they can be modeled independently in each frame using a simple decoder-only (LM-style) transformer architecture.
Through comprehensive evaluations, we show that using dMel allows us to use a single decoder-only model, and achieve high performance on both automatic speech recognition (ASR) and text-to-speech (TTS) tasks. The ASR task validates that dMel preserves semantic information, while the TTS task shows that dMel are useful for high-fidelity acoustic reconstruction of speech. We also compare dMel to other tokenization methods and find that dMel achieves the best WER for ASR task, which indicates that the semantic information is well preserved. Also, dMel achieves the lower WER score for TTS task when using WhisperX [4] for automatic evaluation and we find model trained with dMel can generate long and natural speech samples (see Supplementary Material).
Figure 1: Prior works on speech tokenization use either heavy self-supervised pretrained encoders [2, 16] to extract semantic tokens (and train a separate decoder for it [22]) or learn compression encoder-decoder models with residual vector quantizations [42, 11] to obtain acoustic tokens. By contrast we eliminate the encoder and simply discretize mel-filerbanks (dMel) to encode audio, and use a simple mel-filterbank vocoder [41] to reconstruct speech signals.
@@ 2 Method
In this section, we first introduce our proposed dMel speech tokenization method, which discretizes log mel-filterbanks energies directly into bins. We then describe our unified LM-style transformer model for ASR and TTS tasks, which leverages dMel for speech tokenization. The model architecture is illustrated in Figure 2.
@@
Different from existing VQ-VAE [7, 46, 19, 42] based speech tokenizers, we propose a discretized log mel-filterbanks based speech tokenizer. The outline of the discretization method is shown in Figure 1. Later in the paper, we show that this tokenizer allows the model to process the input speech signal efficiently and capture the relevant acoustic features for both ASR and TTS tasks.
We denote tensors as $\mathbf{X}$ while $\mathbf{X}_{i,...}$ denote the $(i,...)$-th component of tensor $\mathbf{X}$. First, the speech tokenizer takes the input speech signal $\mathbf{x}$ and computes the log mel-filterbanks representation $\mathbf{M}$:
@@
where $\text{Mel}(\cdot)$ represents the function that computes the log mel-filterbanks, $\mathbf{M}\in\mathbb{R}^{T\times N}$, $N$ is the number of log mel-filterbanks and $T$ is the number of frames in the spectrogram.
@@
To discretize the log mel-filterbanks representation $\mathbf{M}$ into speech tokens, we adopt a codebook $\mathbf{C}$. In this paper, we apply a simple linear discretization, so that the codebook $\mathbf{C}\in\mathbb{R}^{2^{K}}$ and its values are evenly spaced in the range of the log mel-filterbanks values:
@@
In practice, we compute the minimum $m$ and maximum $M$ values of log mel-filterbanks across the entire dataset to define the codebook $\mathbf{C}$. Then we map a magnitude $\mathbf{M}_{t,i}$ of every frequency channel $i=1\dots N$ for the time frame $t=1\dots T$ into a bin index of the codebook $\mathbf{C}$ in the following way:
@@
where $\mathbf{S}\in\mathbf{B}^{T\times N}$ represents the discretized log mel-filterbanks (dMel) with $\mathbf{B}=\{j|j=1,2,3,\dots 2^{K}\}$ and $\mathbf{S}_{t}\in\mathbf{B}^{N}$ being the $t$-th speech token. As the codebook $\mathbf{C}$ has $2^{K}$ distinct values and thus number of bins $|\mathbf{B}|=2^{K}$, each speech token is represented by $N\cdot K$ bits where every $K$ bits are used to represent one of $N$ frequency channels.
@@
To reconstruct the speech signal $\mathbf{x}$ from the speech tokens $\mathbf{S}$, we first transform bin indices back to the log mel-filterbanks representation via the codebook $\mathbf{C}$:
@@
Then, we apply a vocoder [41] to transform reconstructed log mel-filterbanks $\hat{\mathbf{M}}_{t,i}$ back into the time domain signal $\mathbf{x}$. The vocoder is trained independently and is not part of the transformer decoder-based model.
@@ 2.2 Unified Speech-Text Transformer Decoder
Figure 2: Unified Speech-Text Transformer Decoder with speech tokens as dMel.
Modeling speech and text sequences jointly is essential for a model to understand and generate both modalities. However, it is challenging to design a unified model that can handle both speech-to-text and text-to-speech effectively. In this work, we apply a unified LM-style transformer model that takes speech and text tokens as input and generates the output tokens in the target sequence. The model is trained in end-to-end on a combined dataset of speech and text pairs, enabling it to learn the joint representations for ASR and TTS tasks. As we show in the rest of the paper, the crucial part for the joint model training is the proper speech tokenization which dMel provides.
@@
$\mathbf{E}_{t}=\text{Linear}(\mathbf{E^{\prime}}_{t}),\,\text{and}\,\,\,\mathbf{E^{\prime}}_{t}=\text{Concatenate}([\text{Embed}(\mathbf{S}_{t,1}),\text{Embed}(\mathbf{S}_{t,2}),\ldots,\text{Embed}(\mathbf{S}_{t,N})]),$ (6)
where $\mathbf{E}_{t}\in\mathbb{R}^{D}$ is the speech token representation. Here, for every time frame $t$, a speech token $\mathbf{S}_{t}$ is processed in parallel and independently for every frequency channel $i$ by $\text{Embed}(\mathbf{S}_{t,i})$ mapping, and then embeddings of all frequency channels are stacked together to form one vector representation $\mathbf{E^{\prime}}_{t}$ for the frame $t$. Finally, the speech token embeddings $\mathbf{E}_{t}$ are fed into the LM-style transformer models for further processing.
We also implemented other popular speech tokenizers including HuBERT-KM [22] and SpeechTokenizer [46] for comparison. The main difference among these speech tokenizers is the codebook size and codes dimension, shown in Table 2. For both HuBERT-KM and SpeechTokenizer the speech tokens are mapped via a learnable linear layer from their dimension to the text embedding dimension $D$ before feeding into the LM-style transformer model.
Table 2: Comparison between different speech tokenizers: dMel (ours), HuBERT-LM and SpeechTokenizer. For dMel we use $N=80$ log mel-filterbanks and $162^{K}=$ 16 values of the codebook $\mathbf{C}$.
@@
To properly model multi-speaker data, we also include speaker embeddings as input to the transformer decoder. The speaker embeddings are extracted from an independent dvector [37] model [^3] We use a pretrained model “Speaker Encoder” from the YourTTS [9] repository https://github.com/Edresson/YourTTS.. We use a learnable linear layer to map the speaker embeddings to the same dimension as the speech and text token embeddings $D$. The speaker representation is optional for ASR task, but required for TTS task. Hence, during the training, it is applied for text-to-speech and ignored for speech-to-text.
@@
To capture the relative distances between tokens in the input sequence, we apply multiplicative relative positional embedding RoPE [35]. This allows the model to learn the positional relationships between speech tokens, text tokens, and speaker embeddings, enhancing its ability to generate coherent output sequences. For positional embeddings we do not distinguish between text, speech and speaker tokens and thus having global positions notation across all of them, see Figure 2.
Robust Training
Compared to LMs, audio frames are highly redundant with strong local correlations. This makes longform generation difficult for models due to exposure bias [6]. To mitigate exposure bias during training, we apply span-masking [29] to the speech token context, masking out multiple random spans of speech frames. The model is trained to predict the next token based on the masked context. This context-masking strategy helps the model learn to generate accurate speech tokens in the presence of missing information, improving its robustness and generalization. It forces the model to attend to the text rather than copying previously inferred speech tokens due to learnt correlations. We also find that span-masking text tokens improves the ASR task.
@@ 3 Experiments
In this section, we begin by evaluating different speech tokenizers through a common practice in the literature: tokenizing speech into discrete units and then reconstructing the speech to assess the quality of the reconstruction. This approach helps gauge the effectiveness of various tokenization techniques. Following this, we present both TTS and ASR results using an LM-style (decoder-only) model with different speech tokens. While most related work focuses solely on speech synthesis, our study encompasses both speech generation and recognition, providing a more comprehensive evaluation of the tokenization methods. We evaluate the performance of our model mainly on the LibriSpeech dataset and compare it with state-of-the-art speech tokenizers, ASR and TTS models.
@@ 3.1 Training Data
We use several open-sourced datasets with paired speech and text transcription to conduct experiments: i) LibriSpeech [25] dataset (CC BY 4.0) consists of English speech recordings (960h, 16kHz) from various speakers ($\sim$2k) and conditions; ii) LibriTTS [43] (CC BY 4.0) dataset (500h) derived from LibriSpeech improves on it with the proper sentence split, text normalization and keeping samples 24kHz; iii) VCTK [40] contains 44h of English speech (108 speakers); iv) LJSpeech [17] (public domain in US) is a single speaker English audio recordings of 16kHz with read speech from LibriVox [^4] https://librivox.org/pages/public-domain/.. While LibriSpeech is used to train ASR and TTS models, LibriTTS, VCTK and LJSpeech are only used to train the TTS.